OpenStudy (anonymous):

Which of these is equal to sin(70 degrees)

OpenStudy (danjs):

recall that the sin of a value, is the y value of the point on the unit circle

OpenStudy (danjs):

so for 70 degrrees (first quadrant) you know that the sin(70) will be a positive number

OpenStudy (danjs):



OpenStudy (danjs):

also: From symmetry Sin(70) = Cos(90 - 70) = cos(20)

OpenStudy (danjs):

and sin(180 - 70) = sin(70)